OpenStudy (anonymous):

Find the effective rate of interest for 6% compounded quarterly.

OpenStudy (anonymous):

helppp!!!

OpenStudy (anonymous):

formula : \[A= P ( 1+\frac{ r }{ n })^{nt}\] r : interest rate as a decimal so 0.06 n: (compounded) since it is quarterly n= 4 and whatever they give you either to find p = principal or A= amount over period of time

OpenStudy (anonymous):

what's effective rate of interest?

OpenStudy (kropot72):

In this problem the actual rate of interest is 6% per year. However the interest is calculated on a quarterly basis and the interest for each quarter is added to the principal. Therefore the effective rate of interest per year will be greater than 6%. This problem can be solved by letting the principal be $1 and the period be 4 quarters = 1 year.\[A =1(1+\frac{0.06}{4})^{4}\] When you calculate the amount A the effective rate of interest can be easily seen.

OpenStudy (anonymous):

so the answer is 1.06 prcnt

OpenStudy (kropot72):

A = 1.06136. Therefore $1.00 has become $1.06136 after 1 year. So the effective rate of interest is\[r _{effective}=\frac{1.06136-1}{1}\times 100=you\ can\ calculate\]

OpenStudy (anonymous):

so 6.26

OpenStudy (anonymous):

thank youuu

OpenStudy (kropot72):

Nearly right. I get 6.136%.